OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} (coshx)' =\sum_{n=0}^{\infty} sinhx\] how ????

OpenStudy (anonymous):

knowing that coshx=\[\sum_{n=0}^{\infty} x^2 x^n /( 2n)!\]

OpenStudy (anonymous):

dose it make any sense to u m8 ?

OpenStudy (anonymous):

did the derivation and didnt have answer that make any sense

OpenStudy (experimentx):

Jeez ... \( (\cos hx)' = \sin hx \)

OpenStudy (anonymous):

i had 2n x^(2n-1)/2n!

OpenStudy (anonymous):

i am talking about the sum of an infinite series here

OpenStudy (experimentx):

\[ \cos hx = \frac{e^x + e^{-x}}{2}\] \[ \sin hx = \frac{e^x - e^{-x}}{2}\]

OpenStudy (experimentx):

Oh .. sorry ... i thought it was hyperbolic function ... I'm out of my mind today

OpenStudy (experimentx):

is that 'h' or 'n' ??

OpenStudy (anonymous):

the derivation of the sum of the coshx is sinhx its true ,

OpenStudy (anonymous):

it seems that u didnt take the calculus II course yet :P

OpenStudy (anonymous):

ok what is the derivative for x! factorial of x ? ?

OpenStudy (anonymous):

you should be taking the derivative with respect to x, not n. anything with just an n is a constant.

OpenStudy (anonymous):

The only thing you should be applying to the derivative of that power series is the power rule.

OpenStudy (anonymous):

\[\cosh (x) = \sum_{n=0}^\infty \frac{x^{2n}}{(2n)!}\]The derivative of this is:\[\frac{d}{dx}\left(\frac{x^{2n}}{(2n)!}\right)=2n\cdot \frac{x^{2n-1}}{(2n)!}=\frac{x^{2n-1}}{(2n-1)!}\]The infinite summation of what you have left will be sinh(x)

OpenStudy (experimentx):

what's the point of summing if you don't have 'n' ...??